# Provable In-Context Learning of Nonlinear Regression with Transformers

### v1

@@ 5.3 Phase Transition under Sharp $L$-Regime
In the sharp curvature regime where $L=\Omega(\frac{1}{\Delta\delta})$, the update $q_{k}^{(t)}$ increases at a rate of $\Theta(\frac{\eta}{K})$ for all $t\leq T_{f}^{*}$. Let $T_{s}^{1}$ denote the durian of Phase I in this regime. Since this early-stage growth (Phase I) mirrors the dynamics under the flat $L$-regime before $T_{f}^{*}$, we have $T_{s}^{1}=T_{f}^{*}=\Theta(\frac{K\log(K\epsilon^{-1})}{\eta\delta^{2}L^{2}\Delta^{2}})$. However, this initial Phase is insufficient for achieving convergence due to the residual error term proportional to $L\cdot\Delta$ in (11). Consequently, training transitions into a second phase (Phase II), during which both gradient terms $g_{k}^{(t)}$ and $g_{k,k^{\prime}}^{(t)}$ become small. This leads to a slower growth rate of $q_{k}^{(t)}$ compared to the earlier phase. We characterize this slower training phase as follows.
Proposition 4 (Phase II: Slow growth of $q_{k}^{(t)}$ under sharp $L$-regime).
@@
The proof of Proposition˜4 is given in Section˜D.1. To ensure convergence of the prediction loss, the update step size of $q_{k}^{(t)}$ decreases to an order dependent on $\epsilon$ in Phase II. This is because, under the sharp $L$-regime where $\delta^{2}L^{2}\Delta^{2}=\Omega(1)$, the gradients become large, as established in Lemma˜2. This phase transition further supports the convergence guarantee in Theorem˜2, demonstrating that even under sharp curvature, the attention mechanism gradually concentrates on the correct feature vector, ultimately enabling accurate prediction.
@@ 6 Experiments Verification
We adopt the data and task distributions from Section˜3.1. Each data point is sampled from a fixed feature set ${v_{k}\in\mathbb{R}^{d},k=1,\dots,K}$, with each feature $v_{k}$ chosen uniformly at random, i.e., $p_{k}=1/K$. Each task involves learning an exponential function of the form $f(x)=c\cdot\exp(\alpha x+\beta)$, where $c$ is drawn from a continuous uniform distribution with lower bound $L$, satisfying (4). Each prompt consists of $N$ randomly sampled inputs $\{x_{i}\}_{i=1}^{N}$ and corresponding outputs $\{y_{i}\}_{i=1}^{N}=\{f(x_{i})\}_{i=1}^{N}$, along with a query token $x_{\mathrm{query}}$. We set the parameters as follows: $d=15$, $K=4$, $N=100$, $\alpha=1$, $\beta=0.5$, and $\Delta=3$. We generate $M=300$ prompts and train the model for $T=400$ epochs.
[…]
@@
(c) Attention scores ($L_{1}=0.1$).
(d) Attention scores for $L_{1}=1$.
[…]
For flat $L$-regime, as shown in Figure˜1(a) and Figure˜1(c), we observe two distinct training phases. For example, with $L=0.1$, the prediction loss rapidly decreases to around 1.01.01.0 by epoch $t=40$, driven by increasing attention on the target feature shown in Figure˜1(c). Subsequently, it steadily declines to near zero by $t=250$. At the same time, $\mathrm{Attn}_{1}^{(t)}$ approaches 111 under the transformer parameter $\theta$. The convergence time shortens with increasing $L$ due to stronger gradient updates, consistent with Theorem˜1, Proposition˜1, Proposition˜2, and Proposition˜3.
As depicted in Figure˜1(b) and Figure˜1(d), the sharp $L$-regime exhibits different three training phases. Specifically, when $L=1$, the prediction loss drops rapidly to approximately 0.60.60.6 by $t=30$, then decreases steadily to around 0.10.10.1 by $t=110$. After that, it converges slowly toward 0 by $t=200$. The dynamics of the attention scores in Figure˜1(d) exhibit the same three-phase progression. These empirical observations are consistent with our theoretical predictions in Theorem˜2 and Proposition˜4.
7 Conclusions
This paper explored how transformers acquire in-context learning (ICL) abilities in nonlinear regression tasks. We analyzed the dynamics of attention scores during training, showing that attention to relevant features increases rapidly before gradually converging, while attention to irrelevant features decays more slowly with oscillations. We showed that different function Lipschitz constants induce flat and sharp $L$ regimes, shaping the convergence behavior differently. For each regime, we derived convergence time bounds that guarantee near-zero prediction error for a broad class of nonlinear functions. Notably, despite convergence speed depending on task properties, transformers consistently attend to tokens with matching features at convergence, highlighting the robustness of ICL.
@@ 7 Conclusions
[…]
In the appendices, we first present detailed proofs of Lemma˜1 and Lemma˜2 in Appendix˜A and Appendix˜B. For ease of exposition, before establishing the two main theorems, we separately analyze the convergence dynamics under both the flat $L$-regime and the sharp $L$-regime.
Specifically, following the proof of Lemma˜1, we prove Proposition˜2 (see Section˜C.1) and Proposition˜3 (see Section˜C.2) as preliminaries for the proof of Theorem˜1 in the flat $L$-regime (see Section˜C.3).
We then proceed to prove Proposition˜4 (see Section˜D.1) as a preliminary for Theorem˜2 (see Section˜D.2) in the sharp $L$-regime.
Finally, based on Theorem˜1 and Theorem˜2, we prove Proposition˜1 (see Appendix˜E).
Appendix A Proof of Lemma˜1
Recall the definition of the prediction error $\mathcal{L}(\theta)$ in (7):
$\mathcal{L}(\theta)=$ $\frac{1}{2}\sum_{k=1}^{K}\mathbb{E}\left[\mathds{1}\{x_{\mathrm{query}}=v_{k}\}(\hat{y}_{\mathrm{query}}-f(v_{k}))^{2}\right]$
@@ Appendix A Proof of Lemma˜1
Next, under the non-degenerate $L$-Lipschitz condition of the function class (see (4)), we have $\sum_{n\neq k}|f(v_{n})-f(v_{k})|=\mathcal{O}(L\Delta)$. Therefore, the sum can be bounded (order-wise) as
@@
$\mathcal{L}(\theta)=\frac{1}{2}\sum_{k=1}^{K}\mathbb{E}\left[\mathds{1}\{x_{\mathrm{query}}=v_{k}\}(1-\mathrm{Attn}_{k}^{(t)})^{2}\cdot\mathcal{O}(L^{2}\Delta^{2})\right].$
@@
Appendix B Proof of Lemma˜2
Similar to the proof of Lemma˜1 in Appendix˜A, we analyze the gradient term of $q_{k,k^{\prime}}^{(t)}$ below.
First, consider $g_{k}^{(t)}$:
$g_{k}^{(t)}$ $=\mathbb{E}\left[\mathds{1}\{x_{\mathrm{query}}=v_{k}\}\,\mathrm{Attn}_{k}^{(t)}\left(f(v_{k})-\sum_{n\in[K]}\mathrm{Attn}_{n}^{(t)}f(v_{n})\right)^{2}\right]$
@@ Appendix B Proof of Lemma˜2
where the last equality follows from the non-degenerate $L$-Lipschitz condition in (4).
@@
Appendix C Proofs for Flat $L$-Regime
C.1 Proof of Proposition˜2

### v2

@@ 5.3 Phase Transition under Sharp $L$-Regime
In the sharp curvature regime where $L=\Omega(\frac{1}{\Delta\delta})$, the update $q_{k}^{(t)}$ increases at a rate of $\Theta(\frac{\eta}{K})$ for all $t\leq T_{f}^{*}$. Let $T_{s}^{1}$ denote the duration of Phase I in this regime. Since this early-stage growth (Phase I) mirrors the dynamics under the flat $L$-regime before $T_{f}^{*}$, we have $T_{s}^{1}=T_{f}^{*}=\Theta(\frac{K\log(K\epsilon^{-1})}{\eta\delta^{2}L^{2}\Delta^{2}})$. However, this initial Phase is insufficient for achieving convergence due to the residual error term proportional to $L\cdot\Delta$ in Eq.˜11. Consequently, training transitions into a second phase (Phase III), during which both gradient terms $g_{k}^{(t)}$ and $g_{k,k^{\prime}}^{(t)}$ become small. This leads to a slower growth rate of $q_{k}^{(t)}$ compared to the earlier phase. We characterize this slower training phase as follows.
Proposition 4 (Phase III: Slow growth of $q_{k}^{(t)}$ under sharp $L$-regime).
@@
The proof of Proposition˜4 is given in Appendix E.1. To ensure convergence of the prediction loss, the update step size of $q_{k}^{(t)}$ decreases to an order dependent on $\epsilon$ in Phase II. This is because, under the sharp $L$-regime where $\delta^{2}L^{2}\Delta^{2}=\Omega(1)$, the gradients become large, as established in Lemma˜2. This phase transition further supports the convergence guarantee in Theorem˜2, demonstrating that even under sharp curvature, the attention mechanism gradually concentrates on the correct feature vector, ultimately enabling accurate prediction.
@@ 6 Experiments Verification
We adopt the data and task distributions from Section˜3.1. Each data point is sampled from a fixed feature set $v_{k}\in\mathbb{R}^{d},k=1,\dots,K$}, with each feature $v_{k}$ chosen uniformly at random, i.e., $p_{k}=1/K$. Each task involves learning a cosine function of the form $f(x)=\frac{L}{c}\cdot\text{cos}(c\cdot x)$, where $c>0$ is a random constant and $L$ is the Lipschitz constant, satisfying Eq.˜4. Each prompt consists of $N$ randomly sampled inputs $\{x_{i}\}_{i=1}^{N}$ and corresponding outputs $\{y_{i}\}_{i=1}^{N}=\{f(x_{i})\}_{i=1}^{N}$, along with a query token $x_{\mathrm{query}}$. We set the parameters as follows: $d=15$, $K=4$, $N=100$, $c=0.5$, and $\Delta=3$. We generate $M=300$ prompts and train the model for $T=400$ epochs. Appendix A presents additional experiments, including attention map dynamics, robustness check with non-uniform feature frequencies, and polynomial-function tasks.
We analyze a simplified transformer model comprising a single block with one-head self-attention and a feedforward network, incorporating layer normalization and ReLU activation, followed by a linear output layer. Our analysis focuses on two key metrics shown in Figure 1: (1) prediction loss dynamics and (2) attention score evolution, evaluated under flat ($\{0.1,0.2,0.4\}$) and sharp ($\{1.0,1.5,2.0\}$) curvature regimes. The prediction loss is computed as the average squared loss over prompts containing query token $v_{k}$. For attention scores, we track $v_{1}$’s self-attention score $\mathrm{Attn}_{1}^{(t)}$ and other features’ attention scores $\mathrm{Attn}_{k,1}^{(t)}$ ($k\in\{2,3,4\}$) on $v_{1}$ at each epoch.
@@
(c) Attention scores for flat $L=0.1$.
(d) Attention scores for sharp $L=1$.
Figure 1: Training dynamics of prediction losses (top) and attention scores (bottom) for two sets of $L$: flat ($\{0.1,0.2,0.4\}$) and sharp ($\{1.0,1.5,2.0\}$).
For flat $L$-regime, as shown in Figure˜1(a) and Figure˜1(c), we observe two distinct training phases. For example, with $L=0.1$, the prediction loss rapidly decreases to around $1.0$ by epoch $t=40$, driven by increasing attention on the target feature shown in Figure˜1(c). Subsequently, it steadily declines to near zero by $t=250$. At the same time, $\mathrm{Attn}_{1}^{(t)}$ approaches $1$ under the transformer parameter $\theta$. The convergence time shortens with increasing $L$ due to stronger gradient updates, consistent with Theorem˜1, Proposition˜1, Proposition˜2, and Proposition˜3.
As depicted in Figure˜1(b) and Figure˜1(d), the sharp $L$-regime exhibits different three training phases. Specifically, when $L=1$, the prediction loss drops rapidly to approximately $0.6$ by $t=30$, then decreases steadily to around $0.1$ by $t=110$. After that, it converges slowly toward 0 by $t=200$. The dynamics of the attention scores in Figure˜1(d) exhibit the same three-phase progression. These empirical observations are consistent with our theoretical predictions in Theorem˜2 and Proposition˜4.
7 Conclusions and Limitations
We presented provable results showing how transformers can learn a broad family of nonlinear tasks in context, identifying two distinct training regimes governed by task curvature. These findings illuminate the basic mechanisms—attention concentration and curvature-dependent gradient dynamics—that underpin ICL. While our analysis makes simplifying assumptions (single head, balanced prompts), it nonetheless captures the essential inductive biases of transformers. By revealing how Lipschitz continuity and feature separation jointly determine convergence and generalization, our theory offers testable predictions for real-world settings (e.g., effects of task smoothness and feature geometry) and provides a principled starting point for extending formal guarantees to more realistic transformer architectures. Future work can relax our assumptions, explore multi-head and multi-layer settings, and examine how richer token distributions interact with the phase transition we identify.
@@ 7 Conclusions and Limitations
In the appendices, we first present additional experimental results in Appendix A. Then we present detailed proofs of Lemma˜1 and Lemma˜2 in Appendix B and Appendix C. For ease of exposition, before establishing the two main theorems, we separately analyze the convergence dynamics under both the flat $L$-regime and the sharp $L$-regime.
Specifically, following the proof of Lemma˜1, we prove Proposition˜2 (see Appendix D.1) and Proposition˜3 (see Appendix D.2) as preliminaries for the proof of Theorem˜1 in the flat $L$-regime (see Appendix D.3).
We then proceed to prove Proposition˜4 (see Appendix E.1) as a preliminary for Theorem˜2 (see Appendix E.2) in the sharp $L$-regime.
Finally, based on Theorem˜1 and Theorem˜2, we prove Proposition˜1 (see Appendix F).
Appendix A Additional Experimental Results
Building on the prediction loss and attention dynamics shown in Figure˜1, we provide a more detailed analysis of the $4\times 4$ attention maps in this appendix. Figure˜2 and Figure˜3 show the attention patterns for the flat ($L=0.1$) and sharp ($L=1$) loss landscapes, respectively, under the experimental conditions of Section˜6.
Figure 2: Dynamics of attention maps under flat $L=0.1$
Figure 3: Dynamics of attention maps under sharp $L=1$
In the attention maps, each grid position $(i,j)$ corresponds to the attention score $\text{Attn}_{i,j}$, where $i,j\in\{1,2,3,4\}$ are the indices of key and query, respectively. An important property is that the sum of attention scores for each column equals one ($\sum_{i}\text{Attn}_{i,j}=1$), given the same query $v_{j}$. The attention maps show a clear trend: the diagonal entries $(i,i)$ become progressively darker over $t$, indicating that the corresponding self-attention scores $\text{Attn}_{i}$ increase for all $i\in\{1,2,3,4\}$. We observe these scores approaching $1$ before $t=300$ for the flat landscape $L=0.1$ and before $t=200$ for the sharp landscape $L=1$. In contrast, the off-diagonal entries $(i,j)$, where $i\neq j$, become lighter, with $\text{Attn}_{i,j}$ converging toward zero. This behavior supports the theoretical findings presented in Proposition˜2 and Proposition˜3 and is directly reflected in the attention score dynamics plotted in Figure˜1(c) and Figure˜1(d).
(a) Prediction loss for flat $L$-set under polynomial functions.
(b) Prediction loss for sharp $L$-set under polynomial functions.
Figure 4: Training dynamics of prediction losses for two sets of $L$: flat ($\{0.1,0.2,0.4\}$) and sharp ($\{1.0,1.5,2.0\}$) under polynomial functions.
To further validate the generality of our theoretical results, we conducted an additional experiment on a different class of nonlinear functions: random second-degree polynomials of the form $f(x)=x^{\top}Ax+b^{\top}x$, where the matrix $A$ and vector $b$ were randomly generated for each task instance. The elements of $A$ and $b$ satisfy $A_{ij},b_{i}\sim\mathcal{N}(0,(\frac{L}{d})^{2})\cdot$. For this experiment, we relaxed the finite feature set assumption by sampling input features from a continuous Gaussian distribution. In other words, there can be infinitely many feature vectors. For the other parameters, we set $d=30,N=150,M=500$. The prediction loss, averaged over $30$ independent runs, is shown below for both the flat $L$-regime and the sharp $L$-regime. The results strongly corroborate our main findings: The flat regime has two training phases, while the sharp regime has three phases. Within both regimes, a larger Lipschitz constant $L$ consistently results in faster convergence of the prediction loss. This provides further evidence that the identified training dynamics and the role of the Lipschitz constant hold for a broader class of nonlinear functions beyond the trigonometric family, and also hold for more general feature set and token sampling process.
Appendix B Proof of Lemma˜1
Recall the definition of the prediction error $\mathcal{L}(P;Q)$ in Eq.˜7:
$\mathcal{L}(P;Q)=$ $\frac{1}{2}\sum_{k=1}^{K}\mathbb{E}\left[\mathds{1}\{x_{\mathrm{query}}=v_{k}\}(\hat{y}_{\mathrm{query}}-f(v_{k}))^{2}\right]$
@@ Appendix B Proof of Lemma˜1
Next, under the non-degenerate $L$-Lipschitz condition of the function class (see Eq.˜4), we have $\sum_{n\neq k}|f(v_{n})-f(v_{k})|=\mathcal{O}(L\Delta)$. Therefore, the sum can be bounded (order-wise) as
@@
$\mathcal{L}(P;Q)=\frac{1}{2}\sum_{k=1}^{K}\mathbb{E}\left[\mathds{1}\{x_{\mathrm{query}}=v_{k}\}(1-\mathrm{Attn}_{k}^{(t)})^{2}\cdot\mathcal{O}(L^{2}\Delta^{2})\right].$
@@
Appendix C Proof of Lemma˜2
To derive $g_{k}$, we first compute the gradient of prediction loss with respect to $Q^{(t)}$ as follows:
$\nabla_{Q^{(t)}}\mathcal{L}=\mathbb{E}\Big[(\hat{y}_{{\text{query}}}-f(v_{k}))^{\top}\frac{\partial\hat{y}_{{\text{query}}}}{\partial Q^{(t)}}\Big]=\mathbb{E}\Big[(\hat{y}_{{\text{query}}}-f(v_{k}))^{\top}\sum_{i\in[N]}\frac{\partial\text{attn}_{i}}{\partial Q^{(t)}}y_{i}\Big].$
We then compute the gradient of $\text{attn}_{i}^{(t)}$ with respect to $Q^{(t)}$:
$\frac{\partial\text{attn}^{(t)}_{i}}{\partial Q^{(t)}}=$ $\frac{e^{{\bar{X}_{i}}^{\top}Q^{(t)}x_{\text{query}}}\cdot{\bar{X}_{i}}^{\top}x_{\text{query}}^{\top}(\sum_{j\in[N]}e^{{\bar{X}_{j}}^{\top}Q^{(t)}x_{\text{query}}})}{(\sum_{j\in[N]}e^{{\bar{X}_{j}}^{\top}Q^{(t)}x_{\text{query}}})^{2}}$
$-\frac{\sum_{j\in[N]}e^{{\bar{X}_{j}}^{\top}Q^{(t)}x_{\text{query}}}\cdot{\bar{X}_{j}}^{\top}x_{\text{query}}^{\top}\cdot e^{{\bar{X}_{i}}^{\top}Q^{(t)}x_{\text{query}}}}{(\sum_{j\in[N]}e^{{\bar{X}_{j}}^{\top}Q^{(t)}x_{\text{query}}})^{2}}$
$=$ $\text{attn}^{(t)}_{i}\cdot{\bar{X}_{i}}^{\top}x_{\text{query}}-\text{attn}^{(t)}_{i}\sum_{j\in[N]}\text{attn}^{(t)}_{j}\cdot\bar{X}_{j}^{\top}x_{\text{query}}$
$=$ $\text{attn}^{(t)}_{i}\sum_{j\in[N]}\text{attn}^{(t)}_{j}(\bar{X}_{i}-\bar{X}_{j}){x_{\text{query}}}^{\top}.$
Substituting into the above expression gives:
$\nabla_{Q^{(t)}}\mathcal{L}=\mathbb{E}\Big[(\hat{y}_{{\text{query}}}^{(t)}-f(v_{k}))\sum_{i,j\in[N]}\text{attn}_{i}^{(t)}\text{attn}^{(t)}_{j}(\bar{X}_{i}-\bar{X}_{j})x_{\text{query}}^{\top}y_{i}\Big].$
For any feature vectors $v_{k}$ and $v_{k^{\prime}}$, we calculate
$v_{k^{\prime}}^{\top}\nabla_{Q^{(t)}}\mathcal{L}\cdot v_{k}=$ $\mathbb{E}\Big[\mathds{1}\{x_{\text{query}}=v_{k}\}(\hat{y}_{\text{query}}-f(v_{k}))\sum_{i,j\in[N]}\text{attn}^{(t)}_{i}\text{attn}_{j}^{(t)}y_{i}v_{k^{\prime}}^{\top}(\bar{X}_{i}-\bar{X}_{j})\Big]$
$=$ $\mathbb{E}\Big[\mathds{1}\{x_{\text{query}}=v_{k}\}(\hat{y}_{\text{query}}-f(v_{k}))\sum_{m,n\in[K]}\sum_{i\in\mathcal{V}_{m}}\sum_{j\in\mathcal{V}_{n}}\text{attn}_{i}^{(t)}\text{attn}_{j}^{(t)}y_{i}v_{k^{\prime}}^{\top}(v_{m}-v_{n})\Big]$
$=$ $\mathbb{E}\Big[\mathds{1}\{x_{\text{query}}=v_{k}\}(\hat{y}_{\text{query}}-f(v_{k}))\sum_{n\in[K]}\sum_{i\in\mathcal{V}_{k^{\prime}}}\sum_{j\in\mathcal{V}_{n}}\text{attn}^{(t)}_{i}\text{attn}^{(t)}_{j}y_{i}v_{k^{\prime}}^{\top}(v_{k^{\prime}}-v_{n})\Big]$
$+\mathbb{E}\Big[\mathds{1}\{x_{\text{query}}=v_{k}\}(\hat{y}_{\text{query}}-f(v_{k}))\sum_{m\in[K]}\sum_{i\in\mathcal{V}_{m}}\sum_{j\in\mathcal{V}_{k^{\prime}}}\text{attn}^{(t)}_{i}\text{attn}^{(t)}_{j}y_{i}v_{k^{\prime}}^{\top}(v_{m}-v_{k^{\prime}})\Big]$
$=$ $\mathbb{E}\Big[\mathds{1}\{x_{\text{query}}=v_{k}\}(\hat{y}_{\text{query}}-f(v_{k}))\text{Attn}^{(t)}_{k^{\prime}}f(v_{k^{\prime}})\sum_{n\in[K]}\text{Attn}^{(t)}_{n}\Big$]
$-\mathbb{E}\Big[\mathds{1}\{x_{\text{query}}=v_{k}\}(\hat{y}_{\text{query}}-f(v_{k}))\text{Attn}^{(t)}_{k^{\prime}}\sum_{m\in[K]}\text{Attn}^{(t)}_{m}f(v_{m})\Big]$
$=$ $\mathbb{E}\Big[\mathds{1}\{x_{\text{query}}=v_{k}\}(\hat{y}_{\text{query}}-f(v_{k}))\text{Attn}^{(t)}_{k^{\prime}}\sum_{m\in[K]}\text{Attn}^{(t)}_{m}(f(v_{k^{\prime}})-f(v_{m}))\Big].$
Since $\hat{y}_{\text{query}}=\sum_{i\in[N]}\text{attn}_{i}y_{i}=\sum_{m\in[K]}\text{Attn}^{(t)}_{m}f(v_{m})$, we obtain
$v_{k^{\prime}}^{\top}\nabla_{Q^{(t)}}\mathcal{L}v_{k}$ $=-\mathbb{E}\Big[\mathds{1}\{x_{\text{query}}=v_{k}\}\text{Attn}^{(t)}_{k^{\prime}}\sum_{n\in[K]}\sum_{m\in[K]}\text{Attn}^{(t)}_{m}\text{Attn}^{(t)}_{n}(f(v_{k})-f(v_{n}))(f(v_{k^{\prime}})-f(v_{m}))\Big]$
$=-\mathbb{E}\Big[\mathds{1}\{x_{\text{query}}=v_{k}\}\text{Attn}^{(t)}_{k^{\prime}}(f(v_{k})-\sum_{n\in[K]}\text{Attn}^{(t)}_{n}f(v_{n}))(f(v_{k^{\prime}})-\sum_{m\in[K]}\text{Attn}^{(t)}_{m}f(v_{m}))\Big].$
We then derive Eq.˜12 by letting $v_{k}=v_{k^{\prime}}$ using this expression:
$g_{k}^{(t)}$ $=-v_{k}^{\top}\nabla_{Q^{(t)}}\mathcal{L}\cdot v_{k}$
$=\mathbb{E}\left[\mathds{1}\{x_{\mathrm{query}}=v_{k}\}\,\mathrm{Attn}_{k}^{(t)}\left(f(v_{k})-\sum_{n\in[K]}\mathrm{Attn}_{n}^{(t)}f(v_{n})\right)^{2}\right]$
@@ Appendix C Proof of Lemma˜2
where the last equality follows from the non-degenerate $L$-Lipschitz condition in Eq.˜4.
@@
Appendix D Proofs for Flat $L$-Regime
D.1 Proof of Proposition˜2